OpenStudy (anonymous):

what is y = 8x - 71 in standard form?

OpenStudy (welshfella):

standard form is ax + by = c where a , b and c are integers

OpenStudy (welshfella):

so you need to rearrange you equation accordingly

OpenStudy (welshfella):

you can start by subtracting y from both sides y - y = 8x - y - 71 0 = 8x - y - 71 now add 71 to both sides

OpenStudy (welshfella):

can you do that?

OpenStudy (anonymous):

71 = 8x - y

OpenStudy (anonymous):

8x - 71 = y ???

OpenStudy (texaschic101):

Ax + By = C y = 8x - 71 --- subtract 8x from both sides y - 8x = -71 -- re-arrange -8x + y = -71 -- since x cannot be a negative, we will multiply entire equation by -1 8x - y = 71 -- standard form questions ?

OpenStudy (welshfella):

- you were on the way to it 71 = 8x - y - but we put the x and y terms on the left so just swapping them around we get 8x - y = 71